OpenStudy (henrietepurina):

The average annual salary of the employees of a company in the year 2005 was $70 thousand. It increased by the same factor each year and in 2006, the average annual salary was $82 thousand. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)?

OpenStudy (henrietepurina):

@Hero

OpenStudy (anonymous):

let r be the factor f(start + n years) = f(0) f^n f(2006)/f(2005) = r = 82000/70000

OpenStudy (henrietepurina):

ok go on... @douglaswinslowcooper

OpenStudy (anonymous):

We do not see the list of your options for the answers.

OpenStudy (henrietepurina):

ok wait a sec...

OpenStudy (henrietepurina):

f(x) = 70(1.17)x f(x) = 82(1.17)x f(x) = 70(2.2)x f(x) = 82(2.2)x

OpenStudy (henrietepurina):

@douglaswinslowcooper

OpenStudy (anonymous):

r = 82/70 = 1.17 f(x) = 70(1.17)^x would be the relationship

OpenStudy (henrietepurina):

i think its a

OpenStudy (henrietepurina):

thanks!